ADVANCED ANALYSIS Assume that the consumption schedule for a private open economy is such that consumption is:

C = 100+ 0.8Y 

Assume further that planned investment lo and net exports Xn are independent of the level of real GDP and constant at lg = 60 and Xn = 10. Government spending (G) is equal to $0. Recall also that, in equilibrium, the real output produced (y) is equal to aggregate expenditures: 

Y=C+Ig+G+Xn 

nstructions: Round your answers to the nearest whole number. 

a. What is the equilibrium level of income or real GDP for this economy? Equilibrium GDPM) = $D 

b. What happens to equilibrium Y if I changes to 40? Equilibrium GDP (Y)= $ . What does this outcome reveal about the size of the spending multiplier? Spending multiplier =
Answer #1

a) Use the equation and find real GDP at equilibrium

Y = C + I + G + NX

Y = 100 + 0.8Y + 60 + 0 + 10

0.2Y = 170

Y = 170/0.2 = 850

Hence equilibrium real output is $850

b) If Investment is reduced to 40, this is a reduction of 20.

Hence income falls by 20 x 1/(1 - 0.8) = 100.

Hence equilibrium GDP is $750

Spending multiplier = 1/(1-MPC)

= 1/(1-0.8)

= 5.